# Markovian Flow Matching: Accelerating MCMC with Continuous Normalizing Flows

### v1

@@ Abstract
Continuous normalizing flows (CNFs) learn the probability path between a reference and a target density by modeling the vector field generating said path using neural networks. Recently, Lipman et al. [42] introduced a simple and inexpensive method for training CNFs in generative modeling, termed flow matching (FM). In this paper, we re-purpose this method for probabilistic inference by incorporating Markovian sampling methods in evaluating the FM objective and using the learned probability path to improve Monte Carlo sampling. We propose a sequential method, which uses samples from a Markov chain to fix the probability path defining the FM objective. We augment this scheme with an adaptive tempering mechanism that allows the discovery of multiple modes in the target. Under mild assumptions, we establish convergence to a local optimum of the FM objective, discuss improvements in the convergence rate, and illustrate our methods on synthetic and real-world examples.
@@ 1 Introduction
where $\hat{\pi}:\mathbb{R}^{d}\rightarrow\mathbb{R}_{+}$ is a continuously differentiable function which can be evaluated pointwise, and $\smash{Z=\int_{\mathbb{R}^{d}}\hat{\pi}(x)\mathrm{d}x}$ is an unknown normalizing constant. We are interested in generating samples from $\pi$ in order to approximate integrals of the form $\pi[f]=\mathbb{E}_{\pi}[f(x)]$, where $f:\mathbb{R}^{d}\rightarrow\mathbb{R}$.
@@
A standard solution to this problem is Markov chain Monte Carlo (MCMC) [61, 10], which relies on the construction of a Markov process which admits the target $\pi$ as its invariant distribution. One of the most broadly applicable and widely studied MCMC methods is the Metropolis-Hastings (MH) algorithm [29], which proceeds in two steps. First, given a current sample $x$, a new sample is proposed according to some proposal distribution $q(\cdot|x)$. Then, this sample is accepted with probability $\smash{\alpha(x,y)=\min\big{\{}1,\tfrac{\pi(y)q(x|y)}{\pi(x)q(y|x)}\big{\}}}$. This strategy generates a Markov chain with the desired stationary distribution and, under mild conditions on the proposal and the target, also ensures that the Markov chain is ergodic [62]. However, for high-dimensional, multi-modal settings, such methods can easily get stuck in local modes, and suffer from very slow mixing times [e.g., 45].
@@
Naturally, the choice of proposal distribution $q(\cdot|x)$ is critical to ensuring that MH MCMC algorithms explore the target distribution within a reasonable number of iterations. A key goal is to obtain proposal distributions with fast mixing times, which can be applied generically to any target distribution. This is particularly challenging in the face of complex, multi-modal (or metastable) distributions, which commonly arise in applications such as genetics [35], protein folding [38], astrophysics [19], and sensor network localisation [34]. On the one hand, local proposals, such as those employed in the Metropolis-Adjusted Langevin Algorithm (MALA) [63] or Hamiltonian Monte Carlo (HMC) [17, 53] struggle to transition between regions of high-probability, resulting in very long decorrelation times and few effective independent samples [e.g., 46]. On the other hand, global proposal distributions must be very carefully designed in order to avoid high rejection rates, particularly in high dimensions [14, 44].
@@
Another popular approach to sampling is variational inference (VI) [75, 31, 58, 8], which obtains a parametric approximation ${\pi}_{\theta^{*}}(x)\approx\pi(x)$ to the target by minimising the Kullback-Leibler (KL) divergence to the target over a parameterized family of distributions $\mathcal{D}_{\theta}=\{\pi_{\theta}:\theta\in\Theta\}$. State-of-the-art VI methods use normalizing flows (NFs), consisting of a sequence of invertible transformations between a reference and a target distribution, to define a flexible variational family [59]. There has also been growing interest in the use of continuous normalizing flows (CNFs), which define a path between distributions using ordinary differential equations [13, 24, 42]. CNFs avoid the need for strong constraints on the flow but, until recently, have been hampered by expensive maximum likelihood training.
@@
In recent years, several works have sought hybrid methods which utilise NFs to enhance the performance of MCMC algorithms; see [25] for a recent review. For example, NFs have been successfully used to precondition complex Bayesian posteriors, significantly improving the performance of existing MCMC methods [e.g., 56, 30, 36, 65]. The synergy between local MCMC proposals and global, flow-informed proposals has also been explored, leading to enhanced mixing rates and effective estimation of multimodal targets [e.g., 21, 64]. In this paper, we continue this promising line of work.
@@
We introduce a new probabilistic inference scheme which integrates CNFs with MCMC sampling techniques. Our approach utilizes flow matching (FM), a scalable, simulation-free training objective for CNFs recently introduced by [42]. This enables, for the first time, the incorporation of CNFs into an adaptive MCMC algorithm. Concretely, our approach augments a local, gradient-based Markov transition kernel with a non-local, flow-informed transition kernel, defined using a CNF. This CNF, and the corresponding transition kernel, are adapted on-the-fly using samples from the chain, which are used to define the probability path for the FM objective. Our scheme also includes an adaptive tempering mechanism, which is essential for discovering multiple modes in complex target distributions. Under mild assumptions, we establish that the flow-network parameters output by our method converge to a local optimum of the FM objective. We then demonstrate empirically the performance of our approach on several synthetic and real-world examples, illustrating comparable or superior performance to other state-of-the-art sampling methods.
@@ 2 Preliminaries
A continuous normalizing flow (CNF) is a continuous-time generative model which is trained to map samples from a base distribution $p_{0}$ to a given target distribution [13]. Let $v_{t}$ be a time-dependent vector field that runs continuously in the unit interval. This vector field can be used to construct a time-dependent diffeomorphic map called a flow $\phi:[0,1]\times\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$, defined via the ordinary differential equation (ODE):
@@
Given a reference density $p_{0}$, and the flow $\phi$, we can generate a probability density path $p:[0,1]\times\mathbb{R}^{d}\rightarrow\mathbb{R}_{+}$ as the pushforward of $p_{0}$ under $\phi$, viz $p_{t}:=[\phi_{t}]_{\sharp}p_{0}$, for $t\in[0,1]$. This yields, via the instantaneous change–of-variables formula [e.g., 12]
@@
where $x_{s}:=\phi_{s}(x)$, and where $\nabla$ is the divergence operator, i.e. the trace of the Jacobian matrix. In modern applications, the vector field $v_{t}$ is often parameterized using a neural network $v_{t}^{\theta}$, in which case the ODE in (2) is referred to as a neural ODE [13]. In turn, this yields a deep parametric model $\phi_{t}^{\theta}$ for the flow $\phi_{t}$, known as a CNF [24].
@@
We would like to learn a CNF which maps between a given reference density $p_{0}$ and a target density $\pi$. Given samples from the target, one approach is to maximize the log-likelihood $\smash{\mathbb{E}_{x\sim\pi}\left[\log p_{1}^{\theta}(x)\right]}$. In practice, however, maximum likelihood training is very slow as both sampling and likelihood evaluation require multiple network passes to solve the ODE in (2). Flow Matching (FM) Lipman et al. [42] provides an alternative, simulation-free method for training CNFs. Let $p_{t}(x)$ be a target probability density path such that $p_{0}=p$ is a simple reference distribution, and $p_{1}\approx\pi$ is approximately equal to the target distribution. Let $v_{t}(x)$ be a vector field which generates this $p_{t}(x)$. Then the FM objective for the CNF vector field $\smash{v_{t}^{\theta}(x)}$ is defined as
@@
In practice, we do not have direct access to the target vector field, $v_{t}(x)$, and so we cannot minimize (4) directly. However, as shown in Lipman et al. [42, Theorem 2], it is equivalent to minimize the conditional flow-matching (CFM) loss
@@ 3 Markovian Flow Matching
In this section, we present our main contribution, an adaptive MCMC algorithm which combines a non-local, flow-informed transition kernel trained via FM, a local Markov transition kernel, and an adaptive annealing schedule. We begin by describing how CNFs can be used within a MH MCMC algorithm.
@@ 3.1 MCMC with Flow Matching
This corresponds to using a transformation-informed proposal in a Markov transition step, an approach which has been successfully applied using (discrete) normalizing flows [56, 41, 30, 25]. There are various possible choices for the proposal distribution on the reference space (see Appendix A). For example, [20] consider an independent MH (IMH) proposal, where i.i.d. samples are drawn from the reference distribution. Here we focus on a flow-informed random-walk, motivated largely by its superior empirical performance in numerical experiments. This proposal performs particularly well on high-dimensional problems, where overfitting of the CNF can be corrected with stochastic steps, while exacerbated by independent proposals [36]. Concretely, our flow-informed random-walk transition kernel, summarized in Algorithm 2 (see Appendix A), can be written as
@@
Thus far, we have assumed that it is possible to train a CNF which maps samples from the reference distribution $p_{0}$ to (an approximation of) the target distribution $\pi$. Clearly, however, the CFM objective is not immediately applicable in our setting, since we do not have access to samples from the target $\pi$. [68] propose two alternatives in this case: (i) use an importance sampling reweighted objective function, or (ii) use samples from a long-run MCMC algorithm (e.g., MALA) as approximate target samples. Both of these approaches, however, have limitations. The former is unlikely to succeed when the proposal distribution differs significantly from the target distribution, while the latter will only perform well when the chosen MCMC method mixes well. We will adopt a different approach, similar in spirit to other flow-informed MCMC algorithms [21, 64, 32], and the Markovian score climbing algorithm in [51], which updates the parameters of a CNF based on a dynamic estimate of the CFM objective obtained via an adaptive MCMC algorithm.
@@ 3.2 Adaptive MCMC with Flow Matching
and where, as elsewhere, $\smash{\alpha(x,y)=\min\big{\{}1,\tfrac{\pi(y)q(x|y)}{\pi(x)q(y|x)}\big{\}}}$ and $b(x)=\int_{\mathbb{R}^{d}}\alpha(x,y)q(\mathrm{d}y|x)$. In principle, however, other choices such as HMC could also be used. Meanwhile, for the non-local MCMC algorithm, we adopt the flow-informed random walk with non-local Markov kernel $P$ defined in (9). Together, assuming alternate local and non-local steps, these two kernels define a Markov chain with Markov transition kernel $\smash{R:=P\circ Q}$, given explicitly by $\smash{R(x,\mathrm{d}y;\pi,\theta)=\int_{\mathcal{Z}}Q(x,\mathrm{d}z;\pi)P(z,\mathrm{d}y;\pi,\theta)}$. In practice, the balance between local and non-local moves is controlled by the hyperparameter $k_{Q}$, which sets the number of local steps before a global step.
@@
Following each MCMC step, the parameters of the flow-informed Markov transition kernel $\smash{P}$ are updated based on a new estimate of $\smash{\mathcal{J}(\theta;\pi)}$. To be precise, suppose we write $\smash{\mu_{t}:=\mu_{0}R^{k}(\cdot,\cdot;\pi,\theta})$ for the distribution of the Markov chain with kernel $\smash{R(\cdot,\cdot;{\pi,\theta})}$ after $k\in\mathbb{N}$ steps, starting from initialisation $\mu_{0}$, where $\smash{R^{k}=R\circ R\cdots\circ R}$. Our objective function is then given by
@@
For our choice of conditional probability path (i.e., the optimal transport path), we can in fact rewrite this objective as [42, Section 4.1]
$\mathcal{J}(\theta;\mu_{k},\sigma_{\mathrm{min}})=\mathbb{E}_{t\sim\mathcal{U}(0,1)}\mathbb{E}_{x_{1}\sim\mu_{k}}\mathbb{E}_{x_{0}\sim p_{0}}\left[||v_{t}^{\theta}(\phi_{t}(x_{0}|x_{1}))-v_{t}(\phi_{t}(x_{0}|x_{1})|x_{1})||^{2}_{2}\right],$ (14)
@@
$\mathcal{J}(\theta;\{x^{i}(k)\}_{i=1}^{N},\sigma_{\mathrm{min}})$ $=\frac{1}{N}\sum_{i=1}^{N}\|v_{t_{i}}^{\theta}(\phi_{t_{i}}(x^{i}_{0}|x^{i}(k)))-v_{t_{i}}(\phi_{t_{i}}(x^{i}_{0}|x^{i}(k))|x^{i}(k))\|_{2}^{2},$ (15)
@@
For complex (e.g., multimodal) target distributions, it can be challenging to learn a CNF that successfully maps between the reference $p_{0}$ and the target $\pi$. For example, if the locations of the modes of the target are not known a priori, and the MCMC chains are initialized far from one or more of the modes, it is unlikely that the local MCMC kernel, and therefore the trained flow, will ever discover these modes [e.g., 21, Section IV.C]. To alleviate this problem, one approach is to iteratively target a sequence of annealed densities $\smash{\{\eta_{t}(x)\}_{t=0:T}}$, which smoothly interpolate between a simple base distribution $\smash{\eta_{0}(x)}$ (e.g., a standard Gaussian), and the target distribution $\smash{\eta_{T}(x):=\pi(x)}$. This idea is central to other Monte Carlo sampling methods such as Sequential Monte Carlo (SMC) [15] and Annealed Importance Sampling (AIS) [52], as well as sampling methods used in score-based generative modelling [e.g., 66]. In our case, the annealed targets act as intermediary steps within the flow-informed MCMC scheme.
In the setting of Bayesian inference, where the target distribution takes the form $\pi(x)\propto\mathcal{L}(\mathcal{D}|x)\pi_{0}(x)$, for a given likelihood $\mathcal{L}(\mathcal{D}|x)$ and prior $\pi_{0}(x)$, a standard way in which to construct $\smash{\{\eta_{t}(x)\}_{t=0:T}}$ is to use a geometric interpolation from the prior to the posterior [52], viz $\eta_{t}(x)\propto\mathcal{L}(\mathcal{D}|x)^{\beta_{t}}\pi_{0}(x)$, where $\beta_{0:T}$ is a sequence of temperatures satisfying $10=\beta_{0}<\beta_{1}<\dots<\beta_{T}=$ 1. In practice, it can be difficult to choose a good sequence of temperatures. One heuristic for adaptively setting this sequence is the effective sample size (ESS). In particular, by setting the ESS to a user-specified percentage $\alpha$ of the number of particles $N$, the next temperature $\beta_{t}$ in the schedule can be determined through the recursive equation [7]
$\beta_{t}=\inf\bigg{\{}\beta_{t-1}<\beta\leq 1:\frac{\left[\frac{1}{N}\sum_{i=1}^{N}w_{i}^{\beta_{t-1}}(\beta)\right]^{2}}{\frac{1}{N}\sum_{i=1}^{N}w_{i}^{\beta_{t-1}}(\beta)^{2}}=\alpha\bigg{\}},$ (16)
where $w_{i}^{\beta_{t-1}}(\beta)=\mathcal{L}(\mathcal{D}|x)^{\beta}\pi_{0}(x)/\mathcal{L}(\mathcal{D}|x)^{\beta_{t-1}}\pi_{0}(x)=\mathcal{L}(\mathcal{D}|x)^{\beta-\beta_{t-1}}$ are new importance weights given the current temperature $\beta_{t-1}$. In practice, we find that the inclusion of this adaptive tempering scheme is essential in the presence of highly multimodal target distributions, enabling the discovery of modes which are not known a priori.
@@
1:Input: $\pi_{0}$, likelihood $\mathcal{L}$, vector field $v_{t}^{\theta}$, reference $p_{0}$, concentration $\sigma_{min}$, initial parameters $\theta_{0}$, target ESS fraction $\alpha$, iterations $K$, number of particles $N$, local kernel $Q$, flow-informed kernel $P$, MCMC steps per resampling step $k_{Q}$, step sizes $\varepsilon_{1:K}$.
@@
3:Sample $x^{i}_{0}\sim\pi_{0}$ for $i=1,\dots,N$ (initialise samples)
@@
7:     end if
8:     if $k\mod k_{Q}+1=0$ then
9:         $x^{i}_{k}\sim P(x^{i}_{k-1},\cdot;\pi_{0}\mathcal{L}^{\beta_{k}},\theta_{k-1})$ for $i=1,\dots,N$ (flow-informed Markov transition).
10:     else
11:         $x^{i}_{k}\sim Q(x^{i}_{k-1},\cdot;\pi_{0}\mathcal{L}^{\beta_{k}})$ for $i=1,\dots,N$ (local Markov transition).
12:     end if
13:     $\theta_{k}=\theta_{k-1}+\varepsilon_{k}\nabla_{\theta}\mathcal{J}(\theta_{k-1}|\{x^{i}_{k}\}_{i=1}^{N},\sigma_{\mathrm{min}})$ (update flow-network parameters)
14:end for

### v2

@@ Abstract
Continuous normalizing flows (CNFs) learn the probability path between a reference distribution and a target distribution by modeling the vector field generating said path using neural networks. Recently, Lipman et al. [45] introduced a simple and inexpensive method for training CNFs in generative modeling, termed flow matching (FM). In this paper, we repurpose this method for probabilistic inference by incorporating Markovian sampling methods in evaluating the FM objective, and using the learned CNF to improve Monte Carlo sampling. Specifically, we propose an adaptive Markov chain Monte Carlo (MCMC) algorithm, which combines a local Markov transition kernel with a non-local, flow-informed transition kernel, defined using a CNF. This CNF is adapted on-the-fly using samples from the Markov chain, which are used to specify the probability path for the FM objective. Our method also includes an adaptive tempering mechanism that allows the discovery of multiple modes in the target distribution. Under mild assumptions, we establish convergence of our method to a local optimum of the FM objective. We then benchmark our approach on several synthetic and real-world examples, achieving similar performance to other state-of-the-art methods, but often at a significantly lower computational cost.
@@ 1 Introduction
where $\hat{\pi}:\mathbb{R}^{d}\rightarrow\mathbb{R}_{+}$ is a continuously differentiable function which can be evaluated pointwise, and $\smash{Z=\int_{\mathbb{R}^{d}}\hat{\pi}(x)\mathrm{d}x}$ is an unknown normalizing constant. We are interested in generating samples from the target distribution $\pi$ in order to approximate integrals of the form $\pi[f]=\mathbb{E}_{\pi}[f(x)]$, where $f:\mathbb{R}^{d}\rightarrow\mathbb{R}$.
@@
A standard solution to this problem is Markov chain Monte Carlo (MCMC) [64, 12], which relies on the construction of a Markov process which admits the target $\pi$ as its invariant distribution. One of the most broadly applicable and widely studied MCMC methods is the Metropolis-Hastings (MH) algorithm [32], which proceeds in two steps. First, given a current sample $x$, a new sample $y$ is proposed according to some proposal distribution $q(\cdot|x)$. Then, this sample is accepted with probability $\smash{\alpha(x,y)=\min\big{\{}1,\tfrac{\pi(y)q(x|y)}{\pi(x)q(y|x)}\big{\}}}$. This strategy generates a Markov chain with the desired stationary distribution and, under mild conditions on the proposal and the target, also ensures that the Markov chain is ergodic [65]. However, for high-dimensional, multi-modal settings, such methods can easily get stuck in local modes, and suffer from very slow mixing times [e.g., 48].
@@
Naturally, the choice of proposal distribution $q(\cdot|x)$ is critical to ensuring that MH MCMC algorithms explore the target distribution within a reasonable number of iterations. A key goal is to obtain proposal distributions with fast mixing times, which can be applied generically to any target distribution. This is particularly challenging in the face of complex, multi-modal (or metastable) distributions, which commonly arise in applications such as genetics [38], protein folding [41], astrophysics [22], and sensor network localization [37]. On the one hand, local proposals, such as those employed in the Metropolis-Adjusted Langevin Algorithm (MALA) [66] or Hamiltonian Monte Carlo (HMC) [20, 56] struggle to transition between regions of high-probability, resulting in very long decorrelation times and few effective independent samples [e.g., 49]. On the other hand, global proposal distributions must be very carefully designed in order to avoid high rejection rates, particularly in high dimensions [17, 47].
@@
Another popular approach to sampling is variational inference (VI) [79, 34, 61, 10], which obtains a parametric approximation ${\pi}_{\theta^{*}}(x)\approx\pi(x)$ to the target by minimising the Kullback-Leibler (KL) divergence to the target over a parameterized family of distributions $\mathcal{D}_{\theta}=\{\pi_{\theta}:\theta\in\Theta\}$. State-of-the-art VI methods use normalizing flows (NFs), which consist of a sequence of invertible transformations between a reference and a target distribution, to define a flexible variational family [62]. There has also been growing interest in the use of continuous normalizing flows (CNFs), which define a path between distributions using ordinary differential equations [15, 27, 45]. CNFs avoid the need for strong constraints on the flow but, until recently, have been hampered by expensive maximum likelihood training.
@@
In recent years, several works have sought hybrid methods which utilize NFs to enhance the performance of MCMC algorithms; see, e.g., [28] for a recent review. For example, NFs have been successfully used to precondition complex Bayesian posteriors, significantly improving the performance of existing MCMC methods [e.g., 59, 33, 39, 68]. The synergy between local MCMC proposals and global, flow-informed proposals has also been explored, leading to enhanced mixing rates and effective estimation of multimodal targets [e.g., 24, 67].
@@
In this paper, we continue this promising line of work, introducing a new probabilistic inference scheme which integrates CNFs with MCMC sampling techniques. Our approach utilizes flow matching (FM), a scalable, simulation-free training objective for CNFs recently introduced by Lipman et al. [45]. This enables, for the first time, the incorporation of CNFs into an adaptive MCMC algorithm. Concretely, our approach augments a local, gradient-based Markov transition kernel with a non-local, flow-informed transition kernel, defined using a CNF. This CNF, and the corresponding transition kernel, are adapted on-the-fly using samples from the chain, which are used to define the probability path for the FM objective. Our scheme also includes an adaptive tempering mechanism, which is essential for discovering multiple modes in complex target distributions. Under mild assumptions, we establish that the flow-network parameters output by our method converge to a local optimum of the FM objective. We then demonstrate empirically the performance of our approach on several synthetic and real-world examples, illustrating comparable or superior performance to other state-of-the-art sampling methods.
@@ 2 Preliminaries
A continuous normalizing flow (CNF) is a continuous-time generative model which is trained to map samples from a base distribution $p_{0}$ to a given target distribution [15]. Let $v_{t}$ be a time-dependent vector field that runs continuously in the unit interval. Under mild conditions, this vector field can be used to construct a time-dependent diffeomorphic map called a flow $\phi:[0,1]\times\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$, defined via the ordinary differential equation (ODE):
@@
Given a reference density $p_{0}:\mathbb{R}^{d}\rightarrow\mathbb{R}_{+}$, and the flow $\phi$, we can generate a probability density path $p:[0,1]\times\mathbb{R}^{d}\rightarrow\mathbb{R}_{+}$ as the pushforward of $p_{0}$ under $\phi$, viz $p_{t}:=[\phi_{t}]_{\sharp}p_{0}$, for $t\in[0,1]$. This yields, via the instantaneous change-of-variables formula [e.g., 14]
@@
[…]
@@
One would typically like to learn a CNF which maps between a given reference density $p_{0}$ and a target density $\pi$. Given samples from the target, one approach is to maximize the log-likelihood $\smash{\mathbb{E}_{x\sim\pi}\left[\log p_{1}^{\theta}(x)\right]}$. In practice, however, maximum likelihood training is very slow as both sampling and likelihood evaluation require multiple network passes to solve the ODE in (2).
Flow Matching (FM) provides an alternative, simulation-free method for training CNFs [45]. Let $p_{t}(x)$ be a target probability density path such that $p_{0}=p$ is a simple reference distribution, and $p_{1}\approx\pi$ is approximately equal to the target distribution. Let $v_{t}(x)$ be a vector field which generates this $p_{t}(x)$. Then the FM objective for the CNF vector field $\smash{v_{t}^{\theta}(x)}$ is defined as
@@
In practice, we do not have direct access to the target vector field, $v_{t}(x)$, and so we cannot minimize (4) directly. However, as shown in Lipman et al. [45, Theorem 2], it is equivalent to minimize the conditional flow-matching (CFM) loss
@@ 3 Markovian Flow Matching
In this section, we present our main contribution, an adaptive MCMC algorithm which combines a non-local, flow-informed transition kernel trained via FM; a local, gradient-based Markov transition kernel; and an adaptive annealing schedule. We begin by describing how CNFs can be used within a MH MCMC algorithm.
@@ 3.1 MCMC with Flow Matching
This corresponds to using a transformation-informed proposal in a Markov transition step, an approach which has been successfully applied using (discrete) normalizing flows [59, 44, 33, 28].
There are various possible choices for the proposal distribution on the reference space (see Appendix A). For example, [23] consider an independent MH (IMH) proposal, where i.i.d. samples are drawn from the reference distribution. Here we focus on a flow-informed random-walk, motivated largely by its superior empirical performance in numerical experiments. This proposal performs particularly well on high-dimensional problems, where overfitting of the CNF can be corrected with stochastic steps, while exacerbated by independent proposals [39]. Concretely, our flow-informed random-walk transition kernel, summarized in Algorithm 2 (see Appendix A), can be written as
@@
Thus far, we have assumed that it is possible to train a CNF which maps samples from the reference distribution $p_{0}$ to (an approximation of) the target distribution $\pi$. Clearly, however, the CFM objective is not immediately applicable in our setting, since we do not have access to samples from the target $\pi$.
Tong et al. [72] propose two alternatives in this case: (i) use an importance sampling reweighted objective function, or (ii) use samples from a long-run MCMC algorithm (e.g., MALA) as approximate target samples. Both of these approaches, however, have limitations. The former is unlikely to succeed when the proposal distribution differs significantly from the target distribution, while the latter will only perform well when the chosen MCMC method mixes well.
In this paper, we adopt a different approach, updating the parameters of a CNF based on a dynamic estimate of the CFM objective obtained via an adaptive MCMC algorithm. This is similar in spirit to other recent flow-informed MCMC algorithms [24, 67, 35], and the Markovian score climbing algorithm in [54].
@@ 3.2 Adaptive MCMC with Flow Matching
and where, as elsewhere, $\smash{\alpha(x,y)=\min\big{\{}1,\tfrac{\pi(y)q(x|y)}{\pi(x)q(y|x)}\big{\}}}$ and $b(x)=\int_{\mathbb{R}^{d}}\alpha(x,y)q(\mathrm{d}y|x)$. In principle, however, other choices such as HMC could also be used.
Meanwhile, for the non-local MCMC algorithm, we adopt the flow-informed random walk with non-local Markov kernel $P$ defined in (9). Together, assuming alternate local and non-local steps, these two kernels define a Markov chain with Markov transition kernel $\smash{R:=P\circ Q}$, given explicitly by $\smash{R(x,\mathrm{d}y;\pi,\theta)=\int_{\mathcal{Z}}Q(x,\mathrm{d}z;\pi)P(z,\mathrm{d}y;\pi,\theta)}$. In practice, the balance between local and non-local moves is controlled by the hyperparameter $k_{Q}$, which sets the number of local steps before a global step.
@@
Following each MCMC step, the parameters of the flow-informed Markov transition kernel $\smash{P}$ are updated based on a new estimate of $\smash{\mathcal{J}(\theta;\pi)}$. To be precise, suppose we write $\smash{\mu_{t}:=\mu_{0}R^{k}(\cdot,\cdot;\pi,\theta})$ for the distribution of the Markov chain with kernel $\smash{R(\cdot,\cdot;{\pi,\theta})}$ after $k\in\mathbb{N}$ steps, starting from initialization $\mu_{0}$, where $\smash{R^{k}=R\circ R\cdots\circ R}$. Our objective function is then given by
@@
For our choice of conditional probability path (i.e., the optimal transport path), we can in fact rewrite this objective as [45, Section 4.1]
$\mathcal{J}(\theta;\mu_{k},\sigma_{\mathrm{min}})=\mathbb{E}_{t\sim\mathcal{U}(0,1)}\mathbb{E}_{x_{1}\sim\mu_{k}}\mathbb{E}_{x_{0}\sim p_{0}}\left[\left|\left|v_{t}^{\theta}(\phi_{t}(x_{0}|x_{1}))-v_{t}(\phi_{t}(x_{0}|x_{1})|x_{1})\right|\right|^{2}_{2}\right],$ (14)
@@
$\mathcal{J}(\theta;\{x^{i}(k)\}_{i=1}^{N},\sigma_{\mathrm{min}})$ $=\frac{1}{N}\sum_{i=1}^{N}\left|\left|v_{t_{i}}^{\theta}(\phi_{t_{i}}(x^{i}_{0}|x^{i}(k)))-v_{t_{i}}(\phi_{t_{i}}(x^{i}_{0}|x^{i}(k))|x^{i}(k))\right|\right|_{2}^{2},$ (15)
@@
For complex (e.g., multimodal) target distributions, it can be challenging to learn a CNF that successfully maps between the reference $p_{0}$ and the target $\pi$. For example, if the locations of the modes of the target are not known a priori, and the MCMC chains are initialized far from one or more of the modes, it is unlikely that the local MCMC kernel, and therefore the trained flow, will ever discover these modes [e.g., 24, Section IV.C]. To alleviate this problem, one approach is to iteratively target a sequence of annealed densities $\smash{\{\pi_{k}(x)\}_{k=0:K}}$, which smoothly interpolate between a simple base distribution $\smash{\pi_{0}(x)}$ (e.g., a standard Gaussian), and the target distribution $\smash{\pi_{K}(x):=\pi(x)}$. This idea is central to other Monte Carlo sampling methods such as Sequential Monte Carlo (SMC) [18] and Annealed Importance Sampling (AIS) [55], as well as sampling methods used in score-based generative modelling [e.g., 70]. In our case, the annealed targets act as intermediary steps within the flow-informed MCMC scheme.
A standard way in which to construct the sequence $\smash{\{\pi_{k}(x)\}_{k=0:K}}$ is to use a geometric interpolation, defining
$\pi_{k}(x)=\pi_{K}(x)^{\beta_{k}}\pi_{0}(x)^{1-\beta_{k}},$ (16)
where $\beta_{0:K}$ is a sequence of temperatures which satisfies $10=\beta_{0}<\beta_{1}<\dots<\beta_{K}=$ 1 [e.g., 55]. In practice, it can be difficult to choose a good sequence of temperatures that provides a smooth transition between densities. One heuristic for adaptively setting this sequence is based on the effective sample size (ESS). In particular, by setting the ESS to a user-specified percentage $\alpha$ of the number of particles $N$, the next temperature $\beta_{k}$ in the schedule can be determined by solving the recursive equation [9]
$\beta_{k}=\inf\bigg{\{}\beta_{k-1}<\beta\leq 1:\frac{\left[\frac{1}{N}\sum_{i=1}^{N}w_{i}^{\beta_{k-1}}(\beta)\right]^{2}}{\frac{1}{N}\sum_{i=1}^{N}w_{i}^{\beta_{k-1}}(\beta)^{2}}=\alpha\bigg{\}},$ (17)
where $w_{i}^{\beta_{k-1}}(\beta)=\left[\pi_{K}(x^{i})^{\beta}\pi_{0}(x^{i})^{1-\beta}\right]/\left[\pi_{K}(x^{i})^{\beta_{k-1}}\pi_{0}(x^{i})^{1-\beta_{k-1}}\right]=[\pi_{K}(x^{i})/\pi_{0}(x^{i})]^{\beta-\beta_{k-1}}$ are new importance weights given the current temperature $\beta_{k-1}$. In practice, we find that the inclusion of this adaptive tempering scheme is essential in the presence of highly multimodal target distributions, enabling the discovery of modes which are not known a priori.
@@
1:Input: target $\pi$, base $\pi_{0}$, vector field $v_{t}^{\theta}$, reference $p_{0}$, concentration $\sigma_{min}$, initial parameters $\theta_{0}$, target ESS fraction $\alpha$, iterations $K$, number of particles $N$, local kernel $Q$, flow-informed kernel $P$, MCMC steps per resampling step $k_{Q}$, step sizes $\varepsilon_{1:K}$.
@@
3:Sample $x^{i}_{0}\sim\pi_{0}$ for $i=1,\dots,N$ (initialize samples)
@@
7:         $\pi_{k}=$ solve (16) (update annealing density)
8:     end if
9:     if $k\mod k_{Q}+1=0$ then
10:         $x^{i}_{k}\sim P(x^{i}_{k-1},\cdot;\pi_{k},\theta_{k-1})$ for $i=1,\dots,N$ (flow-informed Markov transition).
11:     else
12:         $x^{i}_{k}\sim Q(x^{i}_{k-1},\cdot;\pi_{k})$ for $i=1,\dots,N$ (local Markov transition).
13:     end if
14:     $\theta_{k}=\theta_{k-1}+\varepsilon_{k}\nabla_{\theta}\mathcal{J}(\theta_{k-1}|\{x^{i}_{k}\}_{i=1}^{N},\sigma_{\mathrm{min}})$ (update flow-network parameters)
15:end for